OpenStudy (jordans_openstudy):

What is the point-slope form of a line with slope that contains the point (-2, 1)? A. y + 1 = 4/5 (x - 2) B. y - 1 = 4/5 (x + 2) C. y + 1 = 4/5 (x + 2) D. y - 1 = 4/5 (x - 2)

OpenStudy (issimplcalcus):

Coordinates are given in (x , y), the easiest way to solve this problem is plug in -2 for x and 1 for y and see if both sides are equal.

OpenStudy (issimplcalcus):

For example A: (-2,1) -2 = x , 1 = y y + 1 = 4/5 (x - 2) -2 + 1 = 4/5 (1 - 2) -1 = -4/5 This isnt true so we can eliminate A.

OpenStudy (jordans_openstudy):

sorry it froze on lost connection on here. So how would i know if its true ?

OpenStudy (jordans_openstudy):

oh and it should be with slope 4/5*

OpenStudy (issimplcalcus):

If one side equals the other it passes through the point.

OpenStudy (jordans_openstudy):

im so confused with this

OpenStudy (jordans_openstudy):

@Will.H